OpenStudy (anonymous):

What is the taylor polynomial of f(x) = (e^x)/ x centered at x = 0?

OpenStudy (rational):

HINT \[e^x = 1+x+\frac{x^2}{2!} + \frac{x^3}{3!} + \frac{x^4}{4!}+\cdots \]

OpenStudy (anonymous):

is it to divide all terms by x?

OpenStudy (rational):

I think so.. but it wont be a polynomial anymore because the first term will be 1/x

OpenStudy (anonymous):

i see, thank you!

OpenStudy (rational):

yw :) wolfram gives the same http://www.wolframalpha.com/input/?i=taylor+polynomial+e%5Ex%2Fx+at+x%3D0

OpenStudy (anonymous):

even if the function itself is undefined at 0, the taylor polynomial is defined! wow, thank you

OpenStudy (anonymous):

hey are u done with it.